OpenStudy (kawaiiiimya):

Can someone help and explain how you solve it 1) The football team has 30 more members than the basketball teams. If each group had 10 more members, the ratio of their membership would be 3:2. How many members are in each group? Write your answer as football x basketball y

OpenStudy (kawaiiiimya):

2)On a particular day, the wind added 4 miles per hour to Alfonso's rate when he was cycling with the wind and subtracted 4 miles per hour from his rate on his return trip. Alfonso found that in the same amount of time he could cycle 54 miles with the wind, he could go only 30 miles against the wind. What is his normal bicycling speed with no wind?

OpenStudy (kawaiiiimya):

3) On a particular day, the wind added 4 miles per hour to Jaime's rate when she was rowing with the wind and subtracted 4 miles per hour from her rate on her return trip. Jaime found that in the same amount of time she could row 57 miles with the wind, she could go only 33 miles against the wind. What is her normal rowing speed with no wind?

OpenStudy (evoker):

For the first problem if F are football players and B are basketball players the first statement translatesto F=30+B . The second says if you add 10 to each you get the ratio of 3: 2 or in otherwords (F+10)/(B+10)=3/2 So finally do a substitution to solve.